Solve using algebraic method

z = 3000x + 5000y

2x + y <= 16

x + 2y <= 11

x + 3y <= 15




1
Expert's answer
2021-11-22T17:48:28-0500

Solution:

Algebraic method is graphical method in LPP.

Assume we have to maximize z = 3000x + 5000y

2x + y "\le" 16

x + 2y "\le" 11

x + 3y "\le" 15

Plotting these inequations on graph:



ABCD is the feasible region.

At A(0,5), z = 3000(0) + 5000(5) = 15000

At B(3,4), z = 3000(3) + 5000(4) = 29000

At C(7,2), z = 3000(7) + 5000(2) = 31000 (maximum)

At D(8,0), z = 3000(8) + 5000(0) = 24000

Thus, maximum value is 31000 at x=7, y=2.